The random variable z is the standard normal score. Find the number K if P(-k<z<k)=0.3900
Answer #1

Solution :

Given that,  

Using standard normal table ,

P(-k < Z < k) = 0.3900

P(Z < k) - P(Z < -k) = 0.3900

2P(Z < k) - 1 = 0.3900

2P(Z < k) = 1 + 0.3900 = 1.39

P(Z < k) = 1.39 / 2 = 0.695

P(Z < 0.51) = 0.695

K = 0.51